# Liouville polarization and the rigidity of their Lagrangian skeleta in dimension $4$

### v1

@@ 3.2. Bidiscs
is an extendable Liouville polarization of $D(A)\times D(B)$, and there is a tame Liouville form $\lambda$ whose skeleton is $\Gamma_{1}\times\Gamma_{2}$.
@@
Let ($\bigl{(}\left\{(p_{i},a_{i})\right\}_{i=1,\dots,m},\lambda_{A}\bigr{)}$ and ($\bigl{(}\left\{(q_{j},b_{j})\right\}_{j=1,\dots,n},\lambda_{B}\bigr{)}$ be Liouville polarizations of $D(A)$ and $D(B)$ (so $\sum a_{i}=A$ and $\sum b_{j}=B$) with skeleton $\Gamma_{A}$ and $\Gamma_{B}$, respectively. Let $m^{\prime}\leq m$, $n^{\prime}\leq n$, and assume that for $i\leq m^{\prime}$ and $j\leq n^{\prime}$ we are given decompositions
@@
[…]
@@
    that makes $({\bf\Sigma}^{\prime},\lambda^{\prime})$ a Liouville polarization of $D(A)\times D(B)$
@@
${\mathcal{B}}\Bigl{(}\cup_{i}D_{i}^{A}\times D(B)\bigcup\cup_{j}D(A)\times D_{j}^{B},\lambda^{\prime}\Bigr{)}\,=\,\bigl{(}D(A)\times D(B)\bigr{)}\setminus(\Gamma_{A}\times\Gamma_{B}).$
@@
${\mathcal{B}}\bigl{(}\cup_{i}D_{i}^{A},\lambda_{A}^{\prime}\bigr{)}\,=\,D(A)\backslash\Gamma_{A}.$
@@
${\mathcal{B}}\Bigl{(}\cup_{i}D_{i}^{A}\times D(B)\bigcup\cup_{j}D(A)\times D_{j}^{B},\lambda^{\prime}\Bigr{)}\,=\,\bigl{(}D(A)\times D(B)\bigr{)}\setminus(\Gamma_{A}\times\Gamma_{B}).$
@@
3.3. Surgery on Liouville polarizations
The aim of this section is Proposition 3.8 that allows to glue certain components of a Liouville polarization, to obtain a “smoother” Liouville polarization. We will need the following rather obvious statement.
@@ 3.3. Surgery on Liouville polarizations
Let $\Sigma\subset B^{4}(1)$ be a smooth symplectic curve that near the boundary of $B^{4}(1)$ agrees with a collection of complex lines (1-dimensional complex vector spaces). Let $\mu\in\mathbb{R}$. Then there exists a Liouville form $\lambda$ on $B^{4}(1)\backslash\Sigma$, tame along $\Sigma$, with residue $\mu$. (We do not claim that $(B^{4}(1),\Sigma,\lambda)$ is a Liouville polarization.)
Proof.  Since $\Sigma$ coincides with a union of complex lines near $\partial B^{4}(1)$, compactifying $B^{4}(1)$ into $\operatorname{\mathbb{C}P}^{2}(1)$ provides a smooth symplectic curve $\Sigma^{\prime}$ in $\operatorname{\mathbb{C}P}^{2}$. This is a polarization, so by Proposition 2.5 (i) there exists a Liouville form $\lambda^{\prime}$ on $\operatorname{\mathbb{C}P}^{2}\backslash\Sigma^{\prime}$ that is tame along $\Sigma^{\prime}$. Restricting $\lambda^{\prime}$ to $B^{4}(1)\subset\operatorname{\mathbb{C}P}^{2}(1)$ provides a Liouville form $\lambda$ on $B^{4}(1)\backslash\Sigma$ tame along $\Sigma$. It has a well-defined residue which is a constant $\nu$ because $\Sigma$ is assumed to be smooth. Then the form
@@
Let $(M,\omega,{\bf\Sigma},\lambda)$ be a polarized symplectic manifold, either closed or Liouville polarized. Assume that the weights $\mu_{i},\mu_{j}$ of two components $\Sigma_{i},\Sigma_{j}$ of ${\bf\Sigma}$ coincide. Then for any neighbourhood $U$ of $\Sigma_{i}\cap\Sigma_{j}$ there exist:
@@
    A Liouville form $\lambda^{\prime}$ on $M\backslash\Sigma^{\prime}$ tame along $\Sigma^{\prime}$ with residues $\mu_{k}^{\prime}$ that coincides with $\lambda$ on the complement of $U$ and is such that $(M,\omega,{\bf\Sigma}^{\prime},\lambda^{\prime})$ is again a (Liouville) polarized symplectic manifold.
Proof.  We first consider a model in $\mathbb{R}^{4}$. Let $\Sigma_{1}=\{z_{2}=0\}$ and $\Sigma_{2}=\{z_{1}=0\}$ be the complex coordinate lines in $\mathbb{C}^{2}$. For $\varepsilon>0$ let $\chi\colon[0,2\varepsilon]\to[0,1]$ be a smooth function that is $1$ near 00 and vanishes on $[\varepsilon,2\varepsilon]$. For $\delta>0$ define the smooth cylinder
@@
on $\mathbb{R}^{4}\setminus(\Sigma_{1}\cup\Sigma_{2})$, tame along $\Sigma_{1}\cup\Sigma_{2}$ with residue $-\mu$. Since $Z_{\delta}=\Sigma_{1}\cup\Sigma_{2}$ on $B^{4}(2\varepsilon)\setminus B^{4}(\varepsilon)$, Lemma 3.7 provides a tame Liouville form $\lambda_{\delta}$ on $B^{4}(2\varepsilon)\setminus Z_{\delta}$ with residue $-\mu$. Consider the shell
$V\,=\,\{R\in(\varepsilon,2\varepsilon)\}\,=\,B^{4}(2\varepsilon)\setminus\overline{B}{}^{4}(\varepsilon).$
Then $\lambda_{0}$ and $\lambda_{\delta}$ are Liouville forms on $V\setminus Z_{\delta}$, tame along $Z_{\delta}$, with equal residue $-\mu$. By Proposition 2.5 (ii) there exist a smooth bounded closed $1$-form $\vartheta$ on $V\backslash Z_{\delta}$ such that
@@
Since $H_{1}(V\backslash Z_{\delta})$ is generated by two small loops around $Z_{\delta}$ on which $\vartheta$ integrates to 00, $\vartheta$ is the derivative of a smooth function $f\colon V\backslash Z_{\delta}\to\mathbb{R}$. Since $\vartheta$ is bounded, $f$ extends to a Lipschitz function on $V$. Let now $\rho\colon[\varepsilon,2\varepsilon]\to[0,1]$ be a smooth function that equals $1$ near $\varepsilon$ and vanishes near $2\varepsilon$, and define the smooth Liouville form $\lambda^{\prime}$ on $B^{4}(2\varepsilon)\setminus Z_{\delta}$ by
@@
Since $\rho(R)f$ is a Lipschitz function, smooth on $V\setminus Z_{\delta}$, $\lambda^{\prime}$ is tame on $V\setminus Z_{\delta}$ with residue $-\mu$, and the same holds on $\{R\leq\varepsilon\}$. Moreover, $\lambda^{\prime}=\lambda_{0}$ near the boundary of $B^{4}(2\varepsilon)$. This finishes the construction of the interpolating model $\bigl{(}Z_{\delta}\cap B^{4}(2\varepsilon),\lambda^{\prime}\bigr{)}$.
@@ 4. Symplectic embeddings
In this section we prove the symplectic embedding results Theorems 1, 2, 3, and 6. The proofs are all based on the same principle, stated in Theorem 2.20, which reduces the proofs to finding convenient polarizations. We first prove Theorem 6. While Theorem 1 is a direct corollary, Theorem 2 deals with an unbounded domain, whence the proof needs some adjustments. We finally prove Theorem 3, using explicit smooth polarizations of the ball.
@@ 4.1.1. The monotone case
and let $\lambda$ be the Liouville form on $\bigl{(}D(A)\times D(B)\bigr{)}\backslash\Sigma$ provided by Lemma 3.5 so that $({\bf\Sigma},\lambda)$ is an extendable Liouville polarization of $D(A)\times D(B)$ with skeleton $\Gamma_{1}\times\Gamma_{2}$.
We now construct a Liouville polarization of $\Omega^{\prime}:=D(a+b)\times D(Kb)$ for $K$ large enough, to be fixed later in the proof. We refer to Figure 4.1 for an illustration of the construction. Take a regular grid (a smooth line) in $D(a+b)$ that divides $D(a+b)$ into topological discs of area $a$ and $b$, and take a point $p_{1}^{\prime}$ in the disc of area $a$ and $p_{2}^{\prime}$ in the disc of area $b$. For a large integer $K$ (to be chosen later), take a regular grid in $D(Kb)$ that divides $D(Kb)$ into $K$ topological discs of area $b$, and choose a point $q_{j}^{\prime}$ in each disc. Lemma 3.5 yields a Liouville polarization
($\left(\Bigl{\{}\bigl{(}D(a+b)\times\{q_{j}^{\prime}\},b\bigr{)},\bigl{(}\{p_{1}^{\prime}\}\times D(Kb),a\bigr{)},\bigl{(}\{p_{2}^{\prime}\}\times D(Kb),b\bigr{)}\Bigr{\}},\lambda^{\prime}\right)$
@@
is a Liouville polarization of $\Omega^{\prime}$. Now $\Sigma_{1}^{\prime}$ has area $Kb$, $\Sigma_{2}^{\prime}$ has area $K(a+2b)$, and #$\#(\Sigma_{1}^{\prime}\cap\Sigma_{2}^{\prime})=K$. For $K$ large enough we therefore find $m$ disjoint topological discs $D_{1}^{i}$ of area strictly larger than $B$ in $\Sigma_{1}^{\prime}$ and $n$ disjoint topological discs $D_{2}^{j}$ of area strictly larger than $A$ in $\Sigma_{2}^{\prime}$ such that each $D_{1}^{i}$ intersects each $D_{2}^{j}$ exactly once. Hence there exists a symplectic morphism $\phi\colon\widehat{\bf\Sigma}\to{\bf\Sigma}^{\prime}$. Since the first cohomology of a disc vanishes, $\phi$ is $(\alpha_{{\operatorname{st}}},\alpha_{{\operatorname{st}}})$-exact. Theorem 2.20 now guarantees an exact symplectic embedding of $\bigl{(}D(A)\times D(B)\bigr{)}\setminus\left(\Gamma_{1}\times\Gamma_{2}\right)$ into $D(a+b)\times D(Kb)\subset Z^{4}(a+b)$. $\square$
@@
Figure 4.1. $\Sigma_{1}^{\prime}$ and $\Sigma_{2}^{\prime}$ in $\Omega^{\prime}=D(a+b)\times D(Kb)$ for $K=4$.

### v2

@@ 3.2. Bidiscs
is an extendable Liouville polarisation of $D(A)\times D(B)$, and there is a tame Liouville form $\lambda$ whose skeleton is $\Gamma_{1}\times\Gamma_{2}$.
@@
Let $\bigl(\left\{(p_{i},a_{i})\right\}_{i=1,\dots,m},\lambda_{A}\bigr)$ and $\bigl(\left\{(q_{j},b_{j})\right\}_{j=1,\dots,n},\lambda_{B}\bigr)$ be Liouville polarisations of $D(A)$ and $D(B)$ (so $\sum a_{i}=A$ and $\sum b_{j}=B$) with skeleton $\Gamma_{A}$ and $\Gamma_{B}$, respectively. Let $m^{\prime}\leq m$, $n^{\prime}\leq n$, and assume that for $i\leq m^{\prime}$ and $j\leq n^{\prime}$ we are given decompositions
@@
    $\qquad\quad{\bf\Sigma}^{\prime}$ $\!\!:=\!\!$ $\bigl\{\bigl(p_{i}^{\ell}\times D(B),a_{i}^{\ell}\bigr)_{i\leq m^{\prime}}^{\ell=1,2},\,\bigl(D(A)\times q_{j}^{\ell},b_{j}^{\ell}\bigr)_{j\leq n^{\prime}}^{\ell=1,2},$
@@
    that makes $({\bf\Sigma}^{\prime},\lambda^{\prime})$ a Liouville polarisation of $D(A)\times D(B)$
@@
${\mathcal{B}}\Bigl(\bigcup_{i}D_{i}^{A}\times D(B)\cup\bigcup_{j}D(A)\times D_{j}^{B},\lambda^{\prime}\Bigr)\,=\,\bigl(D(A)\times D(B)\bigr)\setminus(\Gamma_{A}\times\Gamma_{B}).$
@@
${\mathcal{B}}\Bigl(\bigcup_{i}D_{i}^{A},\lambda_{A}^{\prime}\Bigr)\,=\,D(A)\backslash\Gamma_{A}.$
@@
${\mathcal{B}}\Bigl(\bigcup_{i}D_{i}^{A}\times D(B)\cup\bigcup_{j}D(A)\times D_{j}^{B},\lambda^{\prime}\Bigr)\,=\,\bigl(D(A)\times D(B)\bigr)\setminus(\Gamma_{A}\times\Gamma_{B}).$
@@
3.3. Surgery on Liouville polarisations
The aim of this section is Proposition 3.8 that allows to glue certain components of a Liouville polarisation, to obtain a “smoother” Liouville polarisation. We will need the following rather obvious statement.
@@ 3.3. Surgery on Liouville polarisations
Let $\Sigma\subset B^{4}(1)$ be a smooth symplectic curve that near the boundary of $B^{4}(1)$ agrees with a collection of complex lines (1-dimensional complex vector spaces). Let $\mu\in\mathbb{R}$. Then there exists a Liouville form $\lambda$ on $B^{4}(1)\backslash\Sigma$, tame along $\Sigma$, with residue $\mu$. (We do not claim that $(B^{4}(1),\Sigma,\lambda)$ is a Liouville polarisation.)
Proof.  Since $\Sigma$ coincides with a union of complex lines near $\partial B^{4}(1)$, compactifying $B^{4}(1)$ into $\operatorname{\mathbb{C}P}^{2}(1)$ provides a smooth symplectic curve $\Sigma^{\prime}$ in $\operatorname{\mathbb{C}P}^{2}$. This is a polarisation, so by Proposition 2.5 (i) there exists a Liouville form $\lambda^{\prime}$ on $\operatorname{\mathbb{C}P}^{2}\backslash\Sigma^{\prime}$ that is tame along $\Sigma^{\prime}$. Restricting $\lambda^{\prime}$ to $B^{4}(1)\subset\operatorname{\mathbb{C}P}^{2}(1)$ provides a Liouville form $\lambda$ on $B^{4}(1)\backslash\Sigma$ tame along $\Sigma$. It has a well-defined residue which is a constant $\nu$ because $\Sigma$ is assumed to be smooth. Then the form
@@
Let $(M,\omega,{\bf\Sigma},\lambda)$ be a polarised symplectic manifold, either closed or Liouville polarised. Assume that the weights $\mu_{i},\mu_{j}$ of two components $\Sigma_{i},\Sigma_{j}$ of $\bf\Sigma$} coincide. Then for any neighbourhood $U$ of $\Sigma_{i}\cap\Sigma_{j}$ there exist:
@@
    A Liouville form $\lambda^{\prime}$ on $M\backslash\Sigma^{\prime}$ tame along $\Sigma^{\prime}$ with residues $\mu_{k}^{\prime}$ that coincides with $\lambda$ on the complement of $U$ and is such that $(M,\omega,{\bf\Sigma}^{\prime},\lambda^{\prime})$ is again a (Liouville) polarised symplectic manifold.
Proof.  We first consider a model in $\mathbb{R}^{4}$. Let $\Sigma_{1}=\{z_{2}=0\}$ and $\Sigma_{2}=\{z_{1}=0\}$ be the complex coordinate lines in $\mathbb{C}^{2}$. For $\varepsilon>0$ let $\chi\colon[0,2\varepsilon]\to[0,1]$ be a smooth function that is $1$ near 0 and vanishes on $\varepsilon,2\varepsilon]$. For $\delta>0$ define the smooth cylinder
@@
on $\mathbb{R}^{4}\setminus(\Sigma_{1}\cup\Sigma_{2})$, tame along $\Sigma_{1}\cup\Sigma_{2}$ with residue $\mu$. Since $Z_{\delta}=\Sigma_{1}\cup\Sigma_{2}$ on $B^{4}(2\varepsilon)\setminus B^{4}(\varepsilon)$, Lemma 3.7 provides a tame Liouville form $\lambda_{\delta}$ on $B^{4}(2\varepsilon)\setminus Z_{\delta}$ with residue $\mu$. Consider the shell
[…]
Then $\lambda_{0}$ and $\lambda_{\delta}$ are Liouville forms on $V\setminus Z_{\delta}$, tame along $Z_{\delta}$, with equal residue $\mu$. By Proposition 2.5 (ii) there exist a smooth bounded closed $1$-form $\vartheta$ on $V\backslash Z_{\delta}$ such that
@@
Since $H_{1}(V\backslash Z_{\delta})$ is generated by two small loops around $Z_{\delta}$ on which $\vartheta$ integrates to 0, $\vartheta$ is the derivative of a smooth function $f\colon V\backslash Z_{\delta}\to\mathbb{R}$. Since $\vartheta$ is bounded, $f$ extends to a Lipschitz function on $V$. Let now $\rho\colon[\varepsilon,2\varepsilon]\to[0,1]$ be a smooth function that equals $1$ near $\varepsilon$ and vanishes near $2\varepsilon$, and define the smooth Liouville form $\lambda^{\prime}$ on $B^{4}(2\varepsilon)\setminus Z_{\delta}$ by
@@
Since $\rho(R)f$ is a Lipschitz function, smooth on $V\setminus Z_{\delta}$, $\lambda^{\prime}$ is tame on $V\setminus Z_{\delta}$ with residue $\mu$, and the same holds on $\{R\leq\varepsilon\}$. Moreover, $\lambda^{\prime}=\lambda_{0}$ near the boundary of $B^{4}(2\varepsilon)$. This finishes the construction of the interpolating model $\bigl(Z_{\delta}\cap B^{4}(2\varepsilon),\lambda^{\prime}\bigr)$.
@@ 4. Symplectic embeddings
In this section we prove the symplectic embedding results Theorems 1, 2, 3, and 6. The proofs are all based on the same principle, stated in Theorem 2.20, which reduces the proofs to finding convenient polarisations. We first prove Theorem 6. While Theorem 1 is a direct corollary, Theorem 2 deals with an unbounded domain, whence the proof needs some adjustments. We finally prove Theorem 3, using explicit smooth polarisations of the ball.
4.1. From symplectic morphisms to exact symplectic morphisms.
In this subsection we prove two lemmas that show that the exactness condition is almost for free. The first lemma is useful if the polarisation of the source is smooth.
Lemma 4.1.
Then there exists an $(\alpha,\alpha^{\prime})$-exact symplectic embedding $\varphi\colon\Sigma\to\overset{\raisebox{-0.90417pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip 6.0mu}{\Sigma^{\prime}}$.
Proof.  A basis of $H_{1}(\Sigma,\mathbb{Z})$ is given by oriented closed curves $\gamma_{1},\dots,\gamma_{2g}$ around the $g$ holes and b1b-1 oriented boundary components $\gamma_{2g+1},\dots,\gamma_{2g+b-1}$, see the red curves in Figure 4.1. Since $g+b-1\leq g^{\prime}$, there exists an embedding $f\colon\Sigma\to\overset{\raisebox{-0.90417pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip 6.0mu}{\Sigma^{\prime}}$ as shown in Figure 4.1.
[…]
Figure 4.1. The embedding $f\colon\Sigma\to\overset{\raisebox{-0.90417pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip 6.0mu}{\Sigma^{\prime}}$ (in the figure it is the inclusion).
Notice that $f_{*}\colon H_{1}(\Sigma)\to H_{1}(\Sigma^{\prime},\partial\Sigma^{\prime})$ is injective, over $\mathbb{Z}$ and hence over $\mathbb{R}$. In our context of oriented surfaces, the de Rham isomorphism gives $H_{1}(\Sigma;\mathbb{R})^{*}=H^{1}(\Sigma)$ and $H_{1}(\Sigma^{\prime},\partial\Sigma^{\prime};\mathbb{R})^{*}=H^{1}_{c}(\overset{\raisebox{-0.90417pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip 6.0mu}{\Sigma^{\prime}})$, where both spaces on the right are the de Rham cohomology groups. Thus, $f^{*}\colon H^{1}_{c}(\overset{\raisebox{-0.90417pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip 6.0mu}{\Sigma^{\prime}})\to H^{1}(\Sigma)$ is surjective.
Since the area of $\Sigma^{\prime}$ is larger than the area of $\Sigma$, Moser’s theorem allows to further assume that $f$ is area-preserving: $f^{*}d\alpha^{\prime}=d\alpha$. Write $\gamma_{i}^{\prime}=f(\gamma_{i})$. Since the classes $\gamma_{i}]$ are linearly independent in $H_{1}(\Sigma)$ and since $f^{*}$ is surjective, there exists a compactly supported closed $1$-form $\vartheta$ on $\Sigma^{\prime}$ with periods
see e.g. [22, $\S$ 10.2]. Thus $\phi_{\vartheta}^{1}\circ f$ is an $(\alpha,\alpha^{\prime})$-exact symplectic embedding of $\Sigma$ into $\Sigma^{\prime}$. The proof of the proposition is complete. $\square$
The next lemma deals with the exactness issue for singular polarisations.
Lemma 4.2.
Let $(\Sigma,d\alpha)$ and $(\Sigma^{\prime},d\alpha^{\prime})$ be compact symplectic multi-curves (with or without boundary) and let $\phi\colon\Sigma\to\Sigma^{\prime}$ be a symplectic morphism whose restriction to each component of $\Sigma$ is $(\alpha,\alpha^{\prime})$-exact. Then there exists a Hamiltonian diffeomorphism $f$ of $\Sigma$ with support in an arbitrarily small neighbourhood of $\operatorname{Sing}\Sigma$, that is the identity near $\operatorname{Sing}\Sigma$, and is such that $\phi\circ f$ is an $(\alpha,\alpha^{\prime})$-exact symplectic morphism.
Proof.  We can assume that $\Sigma$ has at least two components, so that each component $\Sigma_{i}$ contains a singular point. We fix in each $\Sigma_{i}$ a singular point $x_{i}$. If $\Sigma_{i}$ contains at least two singular points, we take for each singular point $x\neq x_{i}$ in $\Sigma_{i}$ two small open discs $D_{x}^{i}$ and $D_{x}^{i}{}^{\prime}$ in $\Sigma_{i}$ with $x\in D_{x}^{i}\subset\overline{D_{x}^{i}}\subset D_{x}^{i}{}^{\prime}$, and a smooth path $\gamma_{x}^{i}$ from $x_{i}$ to $x$ in $\Sigma_{i}$. Define the real numbers
$\delta_{i}(x):=\int_{\gamma_{x}^{i}}\phi^{*}\alpha^{\prime}-\alpha,$
which are responsible for the default of exactness of $\phi$. Let now $H_{i}\colon\Sigma_{i}\to\mathbb{R}$ be smooth functions with support in $\bigcup_{x}D_{x}^{i}{}^{\prime}$ that equal $\delta_{i}(x)$ on $D_{x}^{i}$, and let $f\colon\Sigma\to\Sigma$ be the induced Hamiltonan diffeomorphism. It is well-defined, being the identity near $\operatorname{Sing}\Sigma$ (and hence a symplectic morphism from $\Sigma$ to $\Sigma$). Moreover,
$\int_{\gamma_{x}^{i}}(\phi\circ f)^{*}\alpha^{\prime}=\int_{\gamma_{x}^{i}}\phi^{*}\alpha^{\prime}-\delta_{i}(x)=\int_{\gamma_{x}^{i}}\alpha.$
These equalities imply that for each $i$ the $(\phi\circ f)^{*}\alpha^{\prime}$-action and the $\alpha$-action of every curve in $\Sigma_{i}$ joining two singularities are equal. Indeed, for a curve $\gamma$ in $\Sigma_{i}$ from $x$ to $x_{i}$ (where $x=x_{i}$ is allowed), $\gamma\star\gamma_{x}^{i}$ provides a loop in $\Sigma_{i}$, and since $\phi\circ f$ differs from $\phi$ by pre-composition with a Hamiltonian diffeomorphism and $\phi$ is $(\alpha,\alpha^{\prime})$-exact on $\Sigma_{i}$, so is $\phi\circ f$. Hence
$0$ $=$ $\int_{\gamma\star\gamma_{x}^{i}}(\phi\circ f)^{*}\alpha^{\prime}-\alpha$
$=$ $\int_{\gamma}(\phi\circ f)^{*}\alpha^{\prime}-\alpha-\int_{\gamma_{x}^{i}}(\phi\circ f)^{*}\alpha^{\prime}-\alpha$
$=$ $\int_{\gamma}(\phi\circ f)^{*}\alpha^{\prime}-\alpha.$
And if $\gamma$ joins two arbitrary singularities $x$ and $y$ in $\Sigma_{i}$, then $\gamma\star\overline{\gamma_{y}^{i}}\star\gamma_{x}^{i}$ is a loop in $\Sigma_{i}$, so the latter argument applies and shows the desired equality.
We now claim that $\phi\circ f$ is an $(\alpha,\alpha^{\prime})$-exact symplectic morphism between $\Sigma$ and $\Sigma^{\prime}$. Indeed, we have already observed that the special form of this diffeomorphism makes it $(\alpha,\alpha^{\prime})$-exact on each component of $\Sigma$. And if $\gamma$ is a loop in $\Sigma$ that is not contained in a single component of $\Sigma$, then (a time-translation of) $\gamma$ is a concatenation $\gamma_{1}\star\cdots\star\gamma_{k}$, where each $\gamma_{j}$ is a path in a component of $\Sigma$ connecting two singularities. The integrals of $(\phi\circ f)\alpha^{\prime}$ and $\alpha$ agree on each $\gamma_{j}$ by the previous paragraph, and hence also on $\gamma$. $\square$
Remark 4.3.
The above proof shows that if $\phi$ is already known to preserve the action of all loops in $\Sigma$ that pass only through singularities in a subset $\Gamma\subset\operatorname{Sing}\Sigma$, then the Hamiltonian $H$ can be taken with support in an arbitrarily small neighbourhood of $(\operatorname{Sing}\Sigma)\backslash\Gamma$.
@@ 4.2.1. The monotone case
and let $\lambda$ be the Liouville form on $\bigl(D(A)\times D(B)\bigr)\backslash\Sigma$ provided by Lemma 3.5 so that $({\bf\Sigma},\lambda)$ is an extendable Liouville polarisation of $D(A)\times D(B)$ with skeleton $\Gamma_{1}\times\Gamma_{2}$.
We now construct a Liouville polarisation of $\Omega^{\prime}:=D(a+b)\times D(Kb)$ for $K$ large enough, to be fixed later in the proof. We refer to Figure 4.2 for an illustration of the construction. Take a regular grid (a smooth line) in $D(a+b)$ that divides $D(a+b)$ into topological discs of area $a$ and $b$, and take a point $p_{1}^{\prime}$ in the disc of area $a$ and $p_{2}^{\prime}$ in the disc of area $b$. For a large integer $K$ (to be chosen later), take a regular grid in $D(Kb)$ that divides $D(Kb)$ into $K$ topological discs of area $b$, and choose a point $q_{j}^{\prime}$ in each disc. Lemma 3.5 yields a Liouville polarisation
$\left(\Bigl\{\bigl(D(a+b)\times\{q_{j}^{\prime}\},b\bigr),\bigl(\{p_{1}^{\prime}\}\times D(Kb),a\bigr),\bigl(\{p_{2}^{\prime}\}\times D(Kb),b\bigr)\Bigr\},\lambda^{\prime}\right)$
@@
is a Liouville polarisation of $\Omega^{\prime}$. Now $\Sigma_{1}^{\prime}$ has area $Kb$, $\Sigma_{2}^{\prime}$ has area $K(a+2b)$, and $(\Sigma_{1}^{\prime}\cap\Sigma_{2}^{\prime})=K$. For $K$ large enough we therefore find $m$ disjoint topological discs $D_{1}^{i}$ of area strictly larger than $B$ in $\Sigma_{1}^{\prime}$ and $n$ disjoint topological discs $D_{2}^{j}$ of area strictly larger than $A$ in $\Sigma_{2}^{\prime}$ such that each $D_{1}^{i}$ intersects each $D_{2}^{j}$ exactly once. Hence there exists a symplectic morphism $\phi\colon\widehat{\bf\Sigma}\to{\bf\Sigma}^{\prime}$. Since the first cohomology of a disc vanishes, the restriction of $\phi$ to each components of $\widehat{\Sigma}$ is $(\alpha_{{\operatorname{st}}},\alpha_{{\operatorname{st}}})$-exact, so Lemma 4.2 shows that up to precomposing $\phi$ with a Hamiltonian diffeomorphism of $\widehat{\Sigma}$, we can assume that $\phi$ is an $(\alpha_{\operatorname{st}},\alpha_{\operatorname{st}})$-exact symplectic morphism. Theorem 2.20 now guarantees an exact symplectic embedding of $\bigl(D(A)\times D(B)\bigr)\setminus\left(\Gamma_{1}\times\Gamma_{2}\right)$ into $D(a+b)\times D(Kb)\subset Z^{4}(a+b)$. $\square$
@@
Figure 4.2. $\Sigma_{1}^{\prime}$ and $\Sigma_{2}^{\prime}$ in $\Omega^{\prime}=D(a+b)\times D(Kb)$ for $K=3$.